OpenStudy (anonymous):

Solve: (1/3)(6 - 3t) + 18 = 0 a.16/3 b.-16/3 c.-20 d.20

OpenStudy (anonymous):

6-3t=-18*3=-54 3t=60 t=20

OpenStudy (anonymous):

answer d